Prove that of all rectangles with given perimeter, the square has the greatest area.


1
Expert's answer
2021-03-05T07:25:30-0500

Let P be the fixed perimeter of a rectangle with length x and height y


Hence, P = 2x + 2y and Area A = "xy"


Now, P = 2x -2y


y = "\dfrac{P}{2} - x"


Putting value of y in area A


We get,

A = "x(\dfrac{P}{2} - x) = (\dfrac{P}{2})x - x^2"


Let "A = f(x)"


f(x) = "(\dfrac{P}{2})x - x^2"


Differentiating with respect to 'x'


"f'(x) = \dfrac{P}{2} - 2x"


"f'(x) = 0" when "x = \dfrac{P}{4}"


Hence, "x = \dfrac{P}{4}" gives a local maximum value of the area.


Hence, the square of length "\dfrac{P}{4}" has the greatest area.